# WaveletGPT: Wavelet Inspired LLMs

### v2

@@ 3.3 Wavelet Coefficients by Embedding Dimension Coordinates
Figure 3: (Left) Toy example showing how variations in how embeddings move along token dimension, and how we impose multi-rate structure where different embedding dimensions advance at distinct rates while maintaining causality. Latent space now learns at varying rates for each token, with patterns dispersing from dimension 64 to 0. (Right) Validation loss during pre-training on text-8 with learnable multiscale structure. The model achieves comparable performance nearly twice as fast. When trained for the same number of epochs, we get a performance boost akin to adding additional decoder layers. We also demonstrate the architecture’s performance on text-8 with a 32-dim model, matching the speedup similar seenf for a 128-dim model and for a shallower six layers. For the LRA image benchmark, we observe a 10% performance increase without adding extra parameters.
@@ 3.4 Imposing Structure: Toy Example
In Figure 3, we illustrate a toy example of how we impose structure onto decoder Transformer embeddings. The left side shows eight variations along the token dimension, with onset/sudden bursts at token indices 32, 64, etc., decreasing to zero before rising again. As discussed in the introduction, datasets inherently possess a hierarchical structure, which we capture by imposing it on intermediate Transformer embeddings at each layer. In this example, we retain embeddings at the original resolution for half the dimensions (split by a white line). For the other half, we gradually increase the kernel length across the context and compute the average causally. The final embedding dimension averages over the token dimension with a kernel size equal to the context length (zero-padding if necessary). This creates highways, allowing embeddings to move at different rates: the coordinates from $E/2$ to $E$ move at the Transformer’s original speed, while those from 0 to $E/2$ transition from faster to slower movement. This approach enables the attention mechanism to utilize multi-scale features at varying rates across all layers and tokens, as explored in the next section. Further, these multi-scale structure can be made learnable, driven by just the next token prediction.
@@ 4 Experiments
We explain how we incorporated the idea of infusing wavelets into a large language model pre-training. All of the models are trained from scratch, which required substantial computing. The main aim of these experiments is to show how the performance of the models across three modalities improves with/without doing intermediate modifications on embeddings. We also benchmark on LRA tasks.
4.1 Baseline And Training Setup
Our experiments based on the GPT-2 architecture, feature a stack of 10 Transformer decoder layers with a context length of 512, pretrained from scratch. Each modality—text, symbolic music, and raw waveform—shares the same architecture, using an embedding dimension of 128, a feed-forward dimension of 512, and 8 attention heads. We implement a two-layer feed-forward MLP within the Transformer, each layer matching the feed-forward dimension, rather than the single layer typical in Vaswani et al. (2017). The final decoder outputs to a dense layer of 2048 neurons, followed by a layer matching the vocabulary size: 27 for text8, 256 for raw waveform (Goel et al., 2022; Verma, 2022), and 388 for symbolic music. Baseline models consist of standard Transformer decoder blocks without modified embeddings. For our proposed architecture, we retain half of the embedding coordinates and impose either a fixed or learnable multi-scale structure on the other half for all intermediate layers. We do not compare against larger architectures, as this paper focuses on pre-training from scratch. Instead, we present a scaled-down version of GPT-2 suitable for resource-limited academia, evaluating pre-training performance with and without wavelet-inspired blocks. All models were trained from scratch in TensorFlow Abadi et al. (2016) for 25 epochs, starting with a learning rate of 3e-4, decreasing to 1e-5 when loss plateaued. Each model utilized 1M training points, totaling 500 million tokens, randomly cropped from the dataset. The MLP and attention layers used a default dropout rate of 0.1, with no additional regularization. We measured performance using negative log-likelihood loss, as this method improves the core architecture of the transformer-based GPT - helping achieve the objective we want to achieve: predict the next token correctly. Since we are operating on intermediate embeddings, our work can hopefully generalize to setups with structured data similar to text, raw audio, and symbolic music, where one can go from a fine-grained structure to a coarse structure. As shown in Figure 3, we can impose a multi-scale structure that allows the attention mechanism to not only learn dependencies across various embeddings but also inject some information that can capture coarse and fine-grained structure into these embedding coordinates.
@@ 4.1 Baseline And Training Setup
Figure 4: Results for three modalities: natural language, symbolic music, and raw audio. We see that we achieve much faster performance than the baseline, almost twice as fast on shrunk down GPT like architecture. When trained for the same number of epochs, we see a substantial improvement in the pre-training performance, equivalent to a much larger architecture. The black vertical line denotes the epoch at which our architecture achieves the same performance as our baseline architecture.
@@ 4.2 Performance on modalities
We compared the performance of our baseline architecture across three modalities—text, symbolic music, and audio waveform—with and without wavelet-based intermediate operations. Results showed significant performance improvements in all modalities with the same number of training steps. To illustrate, a 0.04 decrease in validation loss is comparable to going from a 16 to a 64-layer model on text-8 dataset (papers-with code, 2024). As shown in Figure 4, our modified GPT architecture achieves this loss nearly twice as quickly in terms of training steps compared to the original model showing GPT-like architecture can indeed take advantage of the structure that we imposed on half of the embedding dimensions. This speedup, i.e., the number of epochs/steps taken to achieve the same performance (SP: same performance epoch) is even smaller for raw audio, due to quasi-stationary nature of audio signals at smaller time scales (20-30 ms for harmonic sounds). For a sampling rate of 16KHz, a context length of 512 would correspond to 32ms, which may be one of the reasons that some of the coordinates nail down the contents of the context in fewer coordinates onto which we impose structure. The convergence is significantly faster for the raw waveform LLM setup, and achieving nearly twice the speed of text-8 and symbolic music. We also compare the absolute clock run times of our modifications in both learnable/non-learnable setups. In Table 1, we report the time taken to complete one epoch relative to our baseline architecture. Our method is computationally inexpensive, as it primarily involves fixed kernel multiplication or learning a single filter convolutional kernel with variable context lengths across different embedding dimensions.
Table 1: Comparison of the negative-log likelihood (NLL) scores for our architecture with three modalities with/without adding wavelet-based hierarchical structure and learnable wavelet transform.
Modality Baseline Proposed SP Epoch SpeedUp Relative GPU Hours
Text-8 0.93 0.92 14.5 epochs 42% 1.013
Raw Audio 1.84 1.70 3.7 epochs 85% 1.042
Symbolic Music 2.08 2.02 13 epochs 48% 1.059
Text-8 (Learnable) 0.93 0.91 12.9 epochs 48.4% 1.094
Wiki-103 (Learnable) 4.11 4.05 9.5 epochs 62% 1.130
@@ 4.3 Similarities and Differences With EMA
We compare with Exponential Moving Averages (EMA) on intermediate signals. Unlike Haar wavelet which takes fixed window weights, which takes the mean of the signal in the window, EMA uses exponential kernel. Let the signal $x_{i}^{l}(t)$, after the $l^{th}$ layer, be of length equal as context length, with $t$ being the token index from 0 to $L$ at embedding dimension $i$. The modified signal $s_{t}$ is:
$s_{0}=x_{i}^{l}(0)\quad s_{t}=\alpha x_{i}^{l}(t)+(1-\alpha)s_{t-1}$
@@
where $\alpha$, the decay factor, satisfies $10<\alpha<$ 1. Unlike an EMA, our method uses a finite kernel, with zero weights outside a specified length, capturing multi-scale information. In text-8 experiments, we applied EMA on half of the embedding dimensions, with $\alpha$ linearly varying between 0 and 1 for dimensions 64 to 128. This under-performed compared to our baseline, with NLL score of 0.94, while our baseline and proposed method achieved scores of 0.93, 0.92, and 0.91 for non-learnable and learnable cases, respectively. Our method provides a simple, signal processing-based scheme, optimizing weights across multiple resolutions driven by next token prediction, outperforming EMA. Depending on $\alpha$, EMA filter produces an exponential kernel, while we maintain a constant kernel or allow weights learned from scratch optimized for the next token prediction. Further, EMA is an Infinite-Impulse Response (IIR) filter, whereas Haar wavelet based kernel is Finite Impulse Response (FIR) filter. Consequently, for each value update, the contributions from previous samples never reach zero. These can accumulate significantly at longer context lengths for certain $\alpha$. The recursive non-learnable nature of EMA IIR filter always ensures some contribution from all embeddings which may explain the performance degradation, whereas our method uses zero weights outside the kernel length, effectively capturing multi-scale information. We explain more in the Appendix.
4.4 Effect of Depth And Model Dimension
We explore two variants of our architecture for experiments on text-8 – i) reducing model dimension from 128 to 32 ii) reduce the number of layers. For the model with dimension 32 for a 10-layer Transformer decoder architecture with eight heads, it still retains faster performance as a baseline, almost twice as fast as seen in Figure 4, and achieves the performance without doing the modification (as seen as baseline) around ten epochs. For the second experiment, we retain the exact architecture as reported in Table 1. We have 6 Transformer Decoder layers, keeping the rest of the parameters the same (feed-forward dimension four times that of the model dimension, eight attention heads) to see the effect of depth. The model, with Haar inspired modifications, similar to Table 1 results continues to get same performance as baseline twice as fast. Both of these experiments are shown in Figure 3.
4.5 Making multi-scale kernels learnable
We allow each of the kernels to be learnable. In the previous section, we defined the shape of the kernel, and computed approximate signals of intermediate layer activations across all layers, with different resolutions occurring at different embedding dimensions to mimic a causal version of wavelet transform. Now we allow each kernel of length $L$ at a particular level to be learnable for computing the approximate signal for various resolutions, a yet another way to compute it. By making the computation of approximate signal learnable, the model is able to learn how to weight every dimension of every decoder layer as opposed to putting a fixed kernel e.g. exponential weighted average. This as can be seen Algorithm 1 only allows 0.02M (20k) extra parameters to our base decoder architecture. This further improves our performance from 42% to 48% faster speedup to get a similar baseline performance, seen in Figure 4, carried out on the text-8. We also benchmark on Wiki-103 to demonstrate that our method works with the GPT-2 tokenizer. As shown in Table 1, we match the performance of a 10-layer architecture at more than twice the speed. In addition to faster convergence, we see a 3.6-point improvement in perplexity scores over the baseline model. While our architecture, with a 512 context length and 128 model dimension, is a simplified version of GPT-2/3, constrained by academic resources, Section 4.4 shows it scales with model size, highlighting its potential for future improvements for decoder only LLM architectures across modalities and datasets.
5 Long Range Arena Benchmarks
We adapt our architecture for Long-Range Arena (LRA) tasks Tay et al. (2021), which test models on long-range prediction across text, images, and mathematical expressions. These tasks evaluate the model’s ability to handle similarity, structure, and reasoning over extended contexts. We focus on transformer-based architectures, as recently reported by Liu et al. (2024), while other variants include state-space and hybrid models or tweaking attention mechanism. For text, we perform binary classification on the IMDb review dataset (Maas et al., 2011) using byte-level data with a context length of 2048 to determine if a movie review is positive or negative. For images, we use CIFAR-10 from the LRA benchmark, classifying sequences of 3072 pixels into one of ten categories. Lastly, we benchmark on Long ListOps, testing the model’s ability to understand hierarchically structured data in extended contexts. As per LRA paper Tay et al. (2021), ”The dataset is comprised of sequences with a hierarchical structure and operators MAX, MEAN, MEDIAN and SUM_MOD that are enclosed by delimiters (brackets). An example (much shorter) sequence is as follows: INPUT: [MAX 4 3 [MIN 2 3] 1 0 [MEDIAN 1 5 8 9, 2]] OUTPUT: 5. In our task, we use a version of ListOps of sequence lengths of up to 2K to test the ability to reason hierarchically while handling long contexts. In the above example, the model needs to access all tokens and model the logical structure of the inputs to make a prediction. The task is a ten-way classification task and is considerably challenging.”
@@ 5 Long Range Arena Benchmarks
Table 2: Performance on LRA tasks (Tay et al. (2020b)) as reported in Liu et al. (2024). Bold indicates the best-performing model, underlined indicates the second best. We use a baseline architecture for all three datasets (Section 5) and modify intermediate embeddings by imposing a hierarchical structure. Non-transformer based, modified attention based or hybrid architectures are not reported.
Transformer Based Attention Models ListOps Text Image
@@
Linear Trans. (Katharopoulos et al., 2020) 16.13 65.90 42.34
@@
Sparse Transformer (Child et al., 2019) 17.07 63.58 44.24
Performer (Choromanski et al., 2021) 18.01 65.40 42.77
Sinkhorn Transformer (Tay et al., 2020a) 33.67 61.20 41.23
@@
FNET (Lee-Thorp et al., 2022) Non-Causal 37.27 56.10 38.07
WavSPA – Ada Transformer (Zhuang et al., 2024) - Non-Causal 55.40 81.60 55.58
Ours (GPT Baseline With Classification Head) 41.65 65.32 49.81
Ours (WaveletGPT With Classification Head) 57.5 66.38 59.81

### v4

@@ 3.1 Connecting wavelets and LLM embeddings
Figure 3: (Left) Toy example showing embeddings before/after imposing multi-rate structure. Different embedding dimensions advance at distinct rates while maintaining causality, as seen from patterns dispersing from dimension 64 to 0. (Right) Validation loss during pre-training on text-8 with learnable multi-scale structure achieving comparable performance nearly twice as fast/performance boost akin to adding additional decoder layers. Our architecture’s performance on text-8 with a 32-dim model matches the speedup similar to that seen for 128-dim and shallower models. LRA image benchmark, a 10% performance increase without adding any parameters
3.2 Wavelet Coefficients by Embedding Coordinates
One option is to compute the approximate signals for each coordinate signal $x_{(i)}^{l}$ across all decoder layers at levels I to IX. For a context length of 512, this would require nine additional signals with resolutions of 512, 256, 128, 64, 32, 16, 8, 4, and 2, significantly increasing complexity and necessitating major modifications to our GPT model. We propose a novel solution to address this. Instead of computing all levels of approximate signals for every intermediate embedding dimension, we parameterize the level by the embedding dimension index. We want to steer the embeddings only a little into the inductive biases we impose to avoid too much tinkering with what they learn. Transformers have been wildly successful without incorporating any inductive biases. Ideally, we want the best of both worlds, nudging intermediate GPT embeddings in only half of the dimensions. We adjust intermediate GPT embeddings in only half the dimensions. Embeddings from 00 to $E/2$ (coordinates 0 to 64 when $E=128$) remain unchanged. For the rest, we apply processing based on their index $i$. If $x^{l}{(i)}$ is an intermediate embedding after the $l^{th}$ decoder layer along the $i^{th}$ dimension, the modified signal $xn^{l}{(i)}$ equals $x^{l}_{(i)}$ for $i\in[0,E/2]$. For $i>E/2$, we impose structure using an approximate signal calculated from wavelet coefficients corresponding to the index $i$. We use a mapping function $f$ that takes coordinate $i$ (ranging from $E/2$ to $E$) and returns the kernel size corresponding to approximation levels from $I$ to $IX$. The linear function gradually increases from level $I$ (kernel size two at $i=E/2$) to level $IX$ (kernel size 512 at $i=E$, or the coarsest representation i.e., a scalar).
Now, let us find out how we compute the modified new signal $xn^{l}_{(i)}$ that replaces the original intermediate Transformer embeddings $x^{l}_{(i)}$. $f(i)$ is the kernel size for the coordinate $i$. The modified signal is either kept the same or modified as $xn_{(i)}^{l}(k)=\frac{1}{f(i)}\sum_{m=k-f(i)}^{k}x_{(i)}^{l}(m)$ as seen in Algorithm 1. For cases where $k-f(i)<0$, we zero-pad the signal to ensure valid average/kernel computation. Specifically, for the Haar wavelet, the modified signal acts as a causal moving average filter with finite length, averaging the embedding signal along the $i^{\text{th}}$ coordinate with a kernel size determined by $f(i)$. This operation does not introduce new parameters or maintain causality in LLMs to prevent future token leakage, as seen in Equation 4. In Algorithm 1, each value of the modified signal at token $k$ is computed using a convolution with a learned kernel $h(.)$ and variable length $f(i)$, parameterized by the embedding coordinate dimension $i$. Each kernel is learned independently for every signal.
@@ 3.3 Imposing Structure: Toy Example
In Figure 3, we illustrate a toy example of how we impose structure onto decoder Transformer embeddings. The left side shows eight variations along the token dimension, with onset/sudden bursts at token indices 32, 64, etc., decreasing to zero before rising again. As discussed in the introduction, datasets inherently possess a hierarchical structure, which we capture by imposing intermediate Transformer embeddings at each layer. In this example, we retain embeddings at the original resolution for half the dimensions (split by a white line). For the other half, we gradually increase the kernel length across the context and compute the average causally. The final embedding dimension averages over the token dimension with a kernel size equal to the context length (zero-padding if necessary). This creates highways, allowing embeddings to move at different rates: the coordinates from $E/2$ to $E$ move at the Transformer’s original speed, while those from 0 to $E/2$ transition from faster to slower movement. This approach enables the attention mechanism to utilize multi-scale features at varying rates across all layers and tokens, as explored in the next section. Further, this multi-scale structure can be made learnable, driven by just the next token prediction.
@@ 4 Experiments
The main aim of these experiments is to show that the pre-training performance of the models across four modalities improves with/without doing intermediate modifications on embeddings inspired by wavelets. We also benchmark on LRA tasks. We propose a shrunk-down GPT baseline architecture that has the same topology. We do not compare against larger architectures, as this paper focuses on pre-training from scratch, and was written in with access to limited computational resources in academia. We evaluate pre-training performance with and without wavelet-inspired blocks. We only report how well our generative model does for pretraining by quantifying the likelihood scores similar to papers such as Mega-Byte (Yu et al., 2023) and Music Transformer (Huang et al., 2019) that only report NLL scores in the entire paper. We also validate our method across various modalities such as text, audio and music. Further, we benchmark it on various input representations such as raw waveform, MIDI tokens, acoustic tokens, text bytes, characters, and BPE tokens, in addition to math expressions. Our experiments, based on the GPT-2 architecture, have 10 Transformer decoder layers with a context length of 512, trained from scratch. Each modality shares the same architecture, using an embedding dimension of 128, a feed-forward dimension of 512, and 8 attention heads. The final decoder outputs a dense layer of 2048 neurons, followed by a layer matching the vocabulary size  [^2] Vocab of 27 for text8, 256 for raw waveform (Goel et al., 2022; Verma, 2022), 388 for symbolic music, and 1024 for ENCODEC speech tokens, 256 for 8-bit raw pixels, 50257 for BPE tokens.
@@
Figure 4: Results for natural language, symbolic music, and raw audio. We perform faster than baseline, almost twice as fast on shrunk-down GPT. We see substantial gains in pre-training performance for the same epochs, equivalent to a much larger architecture. The black vertical line denotes the epoch at which our architecture achieves the same performance as our baseline architecture.
Baseline models consist of standard Transformer decoder blocks without modified embeddings. We retain half  [^3] The choice of half is a hyper-parameter and it is difficult to optimize for every modality and input represenation due to computational resource constraints. If the optimal split is not half, it will only improve the already strong results of the embedding coordinates for our proposed architecture and impose either a fixed or learnable multi-scale structure on the other half for all intermediate layers. All models were trained from scratch in TensorFlow (Abadi et al., 2016) for 25 epochs, starting with a learning rate of 3e-4, decreasing to 1e-5 when loss plateaued. Each model utilized 1M training points, totalling 500 million tokens, randomly cropped from the dataset. We measured performance using negative log-likelihood loss, as this method improves the core architecture of the transformer-based GPT - helping achieve the objective we want to achieve: predict the next token correctly. Since we are operating on intermediate embeddings, our work can hopefully generalize to setups with structured data similar to text, raw audio, and symbolic music, where one can go from a fine-grained structure to a coarse structure. We impose a multi-scale structure that allows the attention mechanism to learn dependencies across embeddings and inject some information that can capture coarse and fine-grained structures into embedding coordinates while maintaining causality.
@@ 4.1 Performance on modalities
We compared the performance of our baseline architecture across three modalities : text, symbolic music, and audio waveform with and without wavelet-based intermediate operations. Results showed significant performance improvements in all modalities with the same number of training steps. To illustrate, a 0.04 decrease in validation loss is comparable to going from a 16 to a 64-layer model on text-8 dataset (papers-with code, 2024). As shown in Figure 4, our modified GPT architecture achieves this loss nearly twice as quickly in training steps as the original model, showing that GPT-like architecture can take advantage of the structure we imposed on half of the embedding dimensions. This speedup, i.e., the number of epochs/steps taken to achieve the same performance (SP: same performance epoch), is even smaller for raw audio due to the quasi-stationary nature of audio signals at smaller time scales (20-30 ms for harmonic sounds). For a sampling rate of 16KHz, a context length of 512 would correspond to 32ms, which may be one of the reasons that some of the coordinates nail down the contents of the context in fewer coordinates onto which we impose structure. The convergence is significantly faster for the raw waveform LLM setup and achieves nearly twice the speed of text-8 and symbolic music. We also ran benchmarks on LibriSpeech corpus about 1000 hours of speech and acoustic tokens, further strengthening our method is generic to handle several types and modalities of tokens. We run our method on audio classification with Audio Transformer over 200 categories of audio for FSD-50K benchmarks (Fonseca et. al, 2020). We get a performance boost and speedups, thereby showcasing the ubiquity of our proposed method for generative modelling and classification as well. We also compare the absolute clock run times of our modifications in both learnable/non-learnable setups. Table 1 reports the time to complete one epoch relative to our baseline architecture. Our method is computationally inexpensive, as it only involves fixed kernel multiplication or learning a single filter convolutional kernel with variable context lengths along different coordinates.
Modality Baseline Proposed SPE SpeedUp Rel. GPU Hrs
Text-8 0.93 0.92 14.5 42% 1.013
Raw Audio 1.84 1.70 3.7 85% 1.042
Symbolic Music 2.08 2.02 13 48% 1.059
Text-8 (L) 0.93 0.91 12.9 48% 1.094
Wiki-103 (L) 4.11 4.05 9.5 62% 1.130
LibriSpeech (L) 2.43 2.40 9.2 63% 1.110
FSD-50K 40.6% 42.8% 32 65% 1.037
Figure 5: Comparison of the negative-log likelihood (NLL) scores for our architecture across three modalities, with and without wavelet-based fixed/learnable (L) structure. (Left) Table shows the NLL scores and speedup, with Same Performance Epoch (SPE) with baseline as 25 epochs, relative GPU hours. (R) FSD-50K Audio Transformer top-5 accuracy results. Vertical green lines indicate the highest accuracy achieved and the point where the same accuracy is reached 60% faster using our proposed method, with no extra parameters.
4.2 Making multi-scale kernels learnable
We allow each kernel to be learnable. In the previous section, we defined the shape of the kernel and computed approximate signals of intermediate layer activations across all layers, with different resolutions occurring at different embedding dimensions to mimic a causal version of the wavelet transform. Now, we allow each kernel of length $L$ at a particular level to be learnable for computing the approximate signal for various resolutions, yet another way to compute it. By making the computation of approximate signal learnable, the model can learn how to weight every decoder layer dimension instead of putting a fixed kernel, e.g. exponentially weighted average. This as can be seen, Algorithm 1 only allows 20k or 0.2 % extra parameters to our base decoder architecture. We run the experiments both for learnable kernels and fixed kernels (three for each case) due to scarcity of resources. Intuitively there will exist an optimal kernel whose shape is best optimized for the next token prediction for a specific modality of interest than simple Haar wavelet. This further improves our performance from 42% to 48% faster speedup to get a similar baseline performance, seen in Figure 4, carried out on the text-8. We also benchmark on Wiki-103 to demonstrate that our method works with the GPT-2 tokenizer giving even larger gains. As shown in Figure 5, we match the performance of a 10-layer architecture at more than twice the speed. In addition to faster convergence, we see a 3.6-point improvement in perplexity scores over the baseline model for Wiki-103. Section 4.4 shows we scale with model size and depth showing promise for improvements in larger LLM architectures.
4.3 Ablation on Depth And Model Dimension
The aim for these experiments was to see if our model scales with depth of the Transformer and the model dimension. We explore two architecture variants on text-8: (i) reducing the model dimension from 128 to 32 and (ii) reducing the number of layers. The model with 32-dimensional, 10 decoder layers (eight heads) achieves baseline performance in around ten epochs and runs nearly twice as fast (Figure 4). For the second experiment, we retain the architecture from Table 1 but reduce the Transformer decoder to six layers while keeping other parameters unchanged (feed-forward dimension four times the model dimension, eight heads). With Haar-inspired modifications, the model matches baseline performance twice as fast, consistent with the results reported in Section 4.1. While it is difficult to scale the architectures beyond a certain depth and model dimension in academic setups: we believe that by seeing the effect of model dimension and depth holding, we are confident that the findings will extrapolate for much larger/deeper models.
4.4 Audio Classification Benchmark
We explore the strength of our method for a typical audio classification on a standard audio classification benchmark FSD-50K. The goal is to identify the sound categories in 1s of audio correctly. We use a transformer-based architecture similar to Audio Transformer as our baseline model. It consists of 128 convolutional filters with a length of 200 learned over 25ms of audio sampled at 16KHz, yielding patches of 400-length audio samples. The convolutional filter output is then max-pooled across the 25ms window to give a single vector of length 128, i.e. the number of convolutional filters being fed to a Transformer stack of 6 layers with model dimension as 64 similar to (Verma & Chafe, 2023; Verma & Berger, 2021). We report on the top-5 % accuracy and relative gain in mAP scores for our proposed architecture with learnable/non-learnable kernels. We see that we get a performance boost of about 2% and a faster convergence of more than 60%, as reported in Figure 5, for the learnable/fixed kernels and baseline.
@@ 4.5 Similarities and Differences With EMA
We compare Exponential Moving Averages (EMA) on intermediate signals. Unlike the Haar wavelet, which takes fixed window weights, which takes the mean of the signal in the window, EMA uses an exponential kernel. Let the signal $x_{i}^{l}(t)$, after the $l^{th}$ layer, be of length equal to context length, with $t$ being the token index from 0 to $L$ at embedding dimension $i$. The modified signal $s_{t}$ is: $s_{0}=x_{i}^{l}(0)\quad s_{t}=\alpha x_{i}^{l}(t)+(1-\alpha)s_{t-1}$ where $\alpha$, the decay factor, satisfies $10<\alpha<$ 1. Unlike an EMA, our method captures multi-scale information using a finite kernel with zero weights outside a specified length. In text-8 experiments, we applied EMA on half of the embedding dimensions, with $\alpha$ linearly varying between 0 and 1 for dimensions 64 to 128 mimicing our approach with a classical approach. This under-performed compared to our baseline, with an NLL score of 0.94, while our baseline and proposed method achieved scores of 0.93, 0.92, and 0.91 for non-learnable and learnable cases, respectively. Our method provides a simple, signal processing-based scheme that optimizes weights across multiple resolutions driven by next-token prediction and outperforms EMA. Depending on $\alpha$, the EMA filter produces an exponential kernel while we maintain a constant kernel or allow weights learned from scratch optimized for the next token prediction. Further, EMA is an Infinite-Impulse Response (IIR) filter, whereas the Haar wavelet-based kernel is a Finite Impulse Response (FIR) filter. Consequently, for each value update, the contributions from previous samples never reach zero. These can accumulate significantly at longer context lengths for certain $\alpha$. The recursive, non-learnable nature of the EMA IIR filter ensures some contribution from all embeddings, which explains performance degradation. In contrast, our method uses zero weights outside the kernel length, capturing multi-scale information.
@@
Table 1: Performance on LRA tasks ((Tay et al., 2020b)) as reported in (Liu et al., 2024). Bold the best-performing model, and underlined indicates the second-best. We use a baseline GPT baseline (Section 5) and modify intermediate embeddings by imposing a hierarchical structure. Non-transformer-based, modified attention-based or hybrid architectures are not reported.
Attention Based Models ListOps Text Image
@@
Linear Trans. (Vyas et al., 2020) 16.13 65.90 42.34
@@
Sparse Trans. (Child et al., 2019) 17.07 63.58 44.24
Performer (Kaiser et al., 2021) 18.01 65.40 42.77
Sinkhorn (Tay et al., 2020a) 33.67 61.20 41.23
@@
Non-Causal
FNET (Lee-Thorp et al., 2022) 37.27 56.10 38.07
WavSPA (Zhuang et al., 2024) 55.40 81.60 55.58
(Ours) GPT Baseline 41.65 65.32 49.81
(Ours) WaveletGPT 57.5 66.38 59.81
5 Long Range Arena Benchmarks